OpenStudy (anonymous):

Evaluate the function for the given value of x. f(-4)= _______. f(x)={3x^2-6, x>-3 {1/2x+5. x <= -3 (less than or equal to)

OpenStudy (anonymous):

is \(-4>3\) or is \(-4\leq -3\) ?

OpenStudy (anonymous):

The second one.

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} 3x^2-6 & \text{if} & x > -4 \\ \frac{1}{2}x+5& \text{if} & x\leq -3 \end{array} \right. \]

OpenStudy (anonymous):

since \(-4\leq -3\) use the bottom formula and compute \[f(-4)=\frac{1}{2}(-4)+5\]

OpenStudy (anonymous):

No the equation on top, instead of -4 its -3

OpenStudy (anonymous):

sorry that was a typo makes no difference though

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} 3x^2-6 & \text{if} & x > -3 \\ \frac{1}{2}x+5& \text{if} & x\leq -3 \end{array} \right.\]

OpenStudy (anonymous):

the part on the right tells you which formula to use

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

So would it be f(-4)=3? Because when I plugged in -4 to the second equation I got 3

OpenStudy (anonymous):

if i i wrote it correctly \(\frac{1}{2}x+5\) then you are right

OpenStudy (anonymous):

Okay cool. :)